OpenStudy (anonymous):

23m + 18 = 12m + 62 M = ? @DJ♬Øϟϟ¥

OpenStudy (anonymous):

Oh i forgot to tag shay :P

OpenStudy (anonymous):

23m+18=12m+62 subtract 18 23m=12m+44 add 12m 11m=44 divide 11 m=4